OpenStudy (anonymous):

The height a ball bounces is less than the height of the previous bounce due to friction. The heights of the bounces form a geometric sequence. Suppose a ball is dropped from one meter and rebounds to 95% of the height of the previous bounce. What is the total distance traveled by the ball when it comes to rest? -Does the problem give you enough info to solve the problem? -How can you write the general term of the sequence? -What formula should you use to calculate the total distance?

OpenStudy (anonymous):

you drop the ball and it goes down one meter lets put that 1 aside for a moment then it goes up \(.95\) and therefore down \(.95\) for a total distance of \(2\times .95^2\)

OpenStudy (anonymous):

ok that was wrong it goes up \(.95\) and down \(.95\) for a total of \(2\times .95\)

OpenStudy (anonymous):

then it goes up \(.95^2\) and down \(.95^2\) for a total of \(2\times .95^2\)

OpenStudy (anonymous):

leaving the 1 out of it, you get \[2\times .95+2\times .95^2+2\times .95^3+2\times .95^4+...\]

OpenStudy (anonymous):

that is the geometric series you need to add up, then add the 1 at the end

OpenStudy (anonymous):

i am still kind of confused. so thats the series. where do i put the 1? and what formula do i use for the total distance?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

leave the one aside, add it at the end it doesn't fit the pattern

OpenStudy (anonymous):

the sum is found by adding a geometric series \[2(.95+.95^2+.95^3+...)=2\times \frac{.95}{1-.95}\]

OpenStudy (anonymous):

i will try it...

OpenStudy (anonymous):

it is actually easier to do it without a calculator \[\frac{.95}{1-.95}=\frac{.95}{.05}=\frac{95}{5}=19\]

OpenStudy (anonymous):

multiply it by 2 and get 38, then add the 1 and get 39

OpenStudy (anonymous):

this still makes no sense. that isnt even the formula the book uses to find the sum. its \[S _{n}=\frac{a _{1}(1-r ^{n})}{ 1-r }\] where a1 is the first term and r is the common ratio and n is the number of terms.

OpenStudy (anonymous):

@timo86m can you help?

OpenStudy (anonymous):

use geometic sequence

OpenStudy (anonymous):

1+r^1+r^2+....+r^n r = .95